# Wildfire Risk Prediction: A Survey of Recent Advances Using Deep Learning Techniques

### v4

@@ 7.3.2 Other Methods
Figure 11: The sample architecture of the separated spatial and temporal modeling method (Marjani et al., 2024b)
@@ 7.4 Spatiotemporal Prediction Methods
Specifically, separated spatial and temporal modeling involves abstracting spatial features into 1D feature vectors using spatial modeling methods, which are then input into temporal prediction models. Coupled spatiotemporal modeling converts the 1D tensors in RNNs into 3D tensors, incorporating spatial dimensions (rows and columns) and using convolutions or graphs (instead of weight matrix multiplication) to determine a grid cell’s future state based on its neighbors’ current and past states. Lastly, spatially enhanced coupled spatiotemporal modeling introduces CNNs or graph-based methods to extract spatial features before the temporal modeling stage.
@@ 7.4.1 Separated Spatial and Temporal Modeling Methods
The approach of separately extracting spatial and temporal features is quite common. The structure of a typical spatiotemporal wildfire risk prediction model is illustrated in the Fig.11. In this model, multiple CNNs are employed to process multidimensional spatial images at each time step, with dimensionality reduction applied before feeding the data into the LSTM. Finally, the output from the LSTM is mapped to output probabilities using a CNN, or directly without it, along with the application of a sigmoid function.
For example, Jin et al. (2020b) proposed enhancing spatial feature extraction by employing CNNs and GCNs to extract spatial information, which is then passed through a feed-forward network (FFN) to unify the data into a 1D vector before being input into a GRU. The output 1D vector is then restored to its original size using deconvolution layers, performing urban fire situation prediction through an end-to-end semantic segmentation approach. Compared to traditional machine learning methods like RF, XGBoost, as well as GRUN and Conv-GRUN, this model showed significant performance improvements.
@@
Zhang et al. (2022) designed and compared four deep neural networks with identical structures but different depths, combining 2D CNN and LSTM models to predict global monthly burned area maps using seven meteorological and two biophysical predictors from the previous 12 months. Each model reduced spatial dimensions through convolution, ReLU, and max-pooling operations and used fully connected layers to flatten the input patches into 1D features before feeding them into LSTM layers. The primary difference among the models was the number of CNN and LSTM layers. Ultimately, the CNN-LSTM network with two convolutional layers and two LSTM layers was selected for its balance of performance and efficiency. The results showed that the combination of convolution and LSTM methods outperformed standalone LSTM and CNN models.
Furthermore, Marjani et al. (2024b) combined CNN with bidirectional LSTM (BiLSTM) for a study in Australia, using wildfire driving factors from the previous four days and wildfire masks to predict next-day wildfire spread. The method first employed four CNN layers to extract spatial features, which were then flattened and input into a three-layer BiLSTM. The model produced binary outputs through dilated convolution, dense layers, and sigmoid activation functions. Comparative experiments demonstrated that this approach outperformed LSTM and Conv-LSTM methods, and the study also confirmed that using longer time steps improved prediction capabilities. In another study, Li et al. (2021b) not only introduced CNN before the LSTM but also incorporated an attention mechanism within the LSTM to assign different weights to the hidden states, reducing information loss and improving prediction accuracy. Applied in Portugal’s Montesinho Natural Park, the model used various meteorological factors from the Fire Weather Index (FWI), historical burned areas, and time and location data to predict future wildfire risk. Comparisons with SVM, XGBoost, neural networks, RNN, and basic LSTM showed that deep learning models significantly outperformed machine learning models, with LSTM surpassing RNN and the CNN-attention LSTM performing best overall.
@@ 7.4.3 Spatially Enhanced Coupled Spatiotemporal Modeling Methods
In addition, there are studies that combine convolutional layers with ConvLSTM to further emphasize spatial feature extraction. Similar to the separated spatial and temporal sequence wildfire risk prediction models, the spatially enhanced coupled spatiotemporal wildfire risk prediction method, as shown in the Fig.12, first extracts spatial features using convolution operations and reduces dimensionality before feeding the data into a ConvLSTM. For the output of the ConvLSTM, either a CNN is further applied followed by sigmoid mapping, or the sigmoid mapping is directly applied to the output.
Figure 12: A typical spatial feature-enhanced coupled spatiotemporal wildfire risk prediction model proposed by He et al. (2024).
@@
Similarly, He et al. (2024) developed a next-day wildfire risk prediction model for eastern China by combining CNN, ConvLSTM, and vision Transformer models, also employing a branch structure. In this model, eight CNN modules were used to model eight meteorological variables, while another CNN module modeled terrain, vegetation, and anthropogenic factors from the last day of each time step. The sequence feature maps generated by the temporal meteorological variables were input into LSTM, and its output was merged with the other variables’ CNN outputs. Channel attention, spatial attention, and CNN were used to extract higher-order features. To emphasize the high-resolution structural features of vegetation and scattered small settlements, Landsat data were used to generate 30-meter monthly high-resolution NDVI products, and vision Transformer was applied to extract these structural features. The flattened outputs from the vision Transformer and CNN were concatenated and classified using an artificial neural network (ANN) binary classifier. Ablation studies demonstrated the effectiveness of incorporating attention mechanisms, high-resolution spatial structure features, and vision Transformer. Sensitivity analysis also confirmed that meteorological factors were the most important among wildfire drivers.
@@
Finally, Chen et al. (2024a) designed an end-to-end encoder-decoder network structure similar to that of Bhowmik et al. (2023) and Yoon and Voulgaris (2022), combining GCN and ConvLSTM. The GCN component was introduced to model the long-distance spatial dependencies required for global wildfire risk prediction, while mitigating the impact of missing data from ocean regions. Explicitly modeling information flow through the graph edges improved the interpretability of the prediction results. Specifically, the Earth’s land surface pixels, as simulated by the JULES-INFERNO model, were treated as graph nodes, and the correlation coefficients between nodes were used to establish the graph’s edges. Spatial dependencies were modeled using a spatial graph convolution model, followed by LSTM for long-term temporal dependencies. The model was trained and validated using 30 years of simulated data from 1961 to 1990 in the JULES-INFERNO dataset, and its performance was compared against LSTM, ConvLSTM, and convolutional autoencoder with LSTM. The proposed model demonstrated superior performance, with ConvLSTM being the best among the comparative models. Additionally, the study used the Louvain method to identify potential wildfire communities, and integrated gradient analysis revealed that short-term (one-month) wildfire risk is primarily driven by lightning, while long-term (11-month) risk is more influenced by temperature, followed by lightning. Node importance analysis also showed that even geographically distant locations could have strong interactions regarding wildfire risk. Similarly, Zhao et al. (2024) aimed to improve model interpretability by combining causal GNN with LSTM, comparing its performance with LSTM and GRU in predicting future.
7.5 Promising Deep Learning New Methods
Currently, the number and variety of features used in most wildfire risk prediction models have reached a plateau. However, due to the nonlinear and complex spatiotemporal relationships between various factors influencing wildfire ignition and spread, the trend in developing deep learning models for wildfire prediction focuses on enhancing their ability to model long-range spatiotemporal semantic dependencies. This approach aims to capture these subtle relationships more effectively. Specifically, there is a progression from time-series prediction to spatiotemporal sequence prediction and further to spatially enhanced spatiotemporal sequence prediction. Efforts are being made to improve the model’s spatiotemporal scale and resolution by incorporating higher-resolution geospatial data over extended time sequences.
With the rapid advancement of computing hardware, particularly GPUs, the scale of input data for deep learning models has continued to expand. However, current wildfire risk prediction models using RNNs, CNNs, Transformers, or GNNs are still limited by their efficiency and capacity to model large spatiotemporal scales, especially when compared to newer models like Mamba (Gu and Dao, 2024). Moreover, many of these models rely on multi-branch networks that use only images or nodes to capture long-range semantic dependencies, neglecting the potential efficiency gains offered by multimodal approaches, such as incorporating textual descriptions. This section will introduce recent advancements in deep learning for image segmentation tasks using multimodal techniques, followed by an overview of the parameter-efficient Mamba model.
@@ 7.5.1 Multimodal Models
Decoder fusion extracts visual and language features separately using different encoders, such as CNNs for visual data and LSTMs for language data, and designs specific fusion modules to combine these features at the decoder stage. For example, (sweere2022deep) developed Bottom-Up Shift and Bidirectional Attentive Refinement modules to progressively fuse the output features of the language and image modules. Similarly, Jain and Gandhi (2022) utilized CNNs to extract multi-stage image features and LSTMs for text feature extraction, integrating these features through a hierarchical cross-modal aggregation module. In another approach, Wang et al. (2022b) designed a knowledge transfer method to adapt image-level knowledge from Contrastive Language-Image Pretraining (CLIP) into RIS for pixel-level segmentation tasks, implementing a Neck module to fuse outputs from both the text and image encoders. Similarly, Yan et al. (2023) introduced a Fusion Neck module to combine image and text encoding information, alongside a multi-query generator module to aid segmentation.
@@
In summary, multimodal tasks offer greater flexibility and intuitive interaction by incorporating natural language to guide the model, compared to unimodal tasks. Multimodal information provides additional contextual data that unimodal models often lack, allowing the language encoder to construct a relation graph for numerical geospatial relations (Yuan et al., 2024). Therefore, multimodal models, especially RIS, hold significant potential in wildfire risk prediction by capturing subtle spatiotemporal semantic dependencies that are difficult to model using only images or geospatial data.
7.5.2 Mamba

### v5

@@ 6.3.2 Other Methods
While deep learning-based wildfire risk prediction methods often overlap in their reliance on similar wildfire driving data, they do not typically distinguish between wildfire spread simulation and broader wildfire risk prediction, as defined in Section 2.2 of this manuscript. Wildfire spread simulation models are generally tailored to specific scenarios, such as predicting the dynamics of fire spread from known or predefined ignition points in localized regions like parks or mountainous areas (Kondylatos et al., 2023; Jiang et al., 2023; Li et al., 2023c; Marjani et al., 2024a). In contrast, wildfire risk prediction models aim to identify areas with potential ignition points or high susceptibility to fire spread, explicitly incorporating the stochastic nature of ignition events. This approach allows for a comprehensive assessment of fire risk by considering both ignition probability and subsequent spread potential. Moreover, wildfire risk prediction accommodates a wider range of spatial scales, from local parks to provincial, national, continental, or even global levels (Santopaolo et al., 2021; Bergado et al., 2021; Chen et al., 2022; Prapas et al., 2023). This adaptability, coupled with the ability to model stochastic ignition events, highlights the broader applicability and versatility of risk prediction models for large-scale wildfire management.
Additionally, the training process for deep learning-based spread simulation models often involves datasets of discrete historical wildfire events combined with associated driving data, which can limit generalization across diverse wildfire scenarios (Huot et al., 2022; Prapas et al., 2022; Kondylatos et al., 2023). Consequently, while spread simulation methods offer precise fire spread modeling for individual events, their predictive performance tends to be less robust when applied to large-scale wildfire risk assessments (Gerard et al., 2023). This distinction underlines the need for improved methodologies explicitly designed for large-scale and long-term wildfire risk prediction, ensuring scalability and reliability across varying geographies and timeframes.
@@ 6.4 Spatiotemporal Prediction Methods
Specifically, separated spatial and temporal modeling involves abstracting spatial features into 1D feature vectors using spatial modeling methods, which are then input into temporal prediction models. Coupled spatiotemporal modeling converts the 1D tensors in RNNs into 3D tensors, incorporating spatial dimensions (rows and columns) and using convolutions or graphs (instead of weight matrix multiplication) to determine a grid cell’s future state based on its neighbors’ current and past states. Lastly, spatially enhanced coupled spatiotemporal modeling introduces CNNs or graph-based methods to extract spatial features before the spatiotemporal modeling stage.
@@ 6.4.1 Separated Spatial and Temporal Modeling Methods
The approach of separately extracting spatial and temporal features is widely adopted in wildfire risk prediction. A typical architecture of such a spatiotemporal model is illustrated in Fig. 5. The model takes as input spatial data composed of multiple driving factors across $n$ time steps ($T_{1}$ through $T_{n}$), represented as tensors with dimensions height ($H$), width ($W$), and number of drivers ($C$). Each temporal slice is processed by parallel spatial feature extraction modules, either with shared or independent weights. The resulting spatial features are then aggregated through a multi-temporal feature fusion stage, during which the original spatial resolution $H\times W$ is typically reduced to a global representation (e.g., $1\times 1$). This fused representation is then passed through a temporal feature extraction module to capture dynamic patterns across the time sequence. Finally, the integrated spatiotemporal features are input into a prediction head to produce the final wildfire risk assessment.
Figure 5: The diagram of the separated spatial and temporal modeling method.
Typically, multiple CNNs are employed to process multidimensional spatial images at each time step, with dimensionality reduction applied before feeding the data into the LSTM. Finally, the output from the LSTM is mapped to output probabilities using a CNN, or directly without it, along with the application of a sigmoid function. This architecture effectively leverages both the spatial pattern recognition capabilities of CNNs and the sequential modeling strengths of LSTMs to capture the complex spatiotemporal dynamics inherent in wildfire systems.
This method is widely used in wildfire risk prediction. For example, Jin et al. (2020b) proposed enhancing spatial feature extraction by employing CNNs and GCNs to extract spatial information, which is then passed through a feed-forward network to unify the data into a 1D vector before being input into a GRU. The output 1D vector is then restored to its original size using deconvolution layers, performing urban fire situation prediction through an end-to-end semantic segmentation approach. Compared to traditional machine learning methods like RF and XGBoost, as well as GRUN and Conv-GRUN, this model showed significant performance improvements.
@@
Zhang et al. (2022) designed and compared four deep neural networks with identical structures but different depths, combining 2D CNN and LSTM models to predict global monthly burned area maps using seven meteorological and two biophysical predictors from the previous 12 months. Each model reduced spatial dimensions through convolution and max-pooling operations and used fully connected layers to flatten the input patches into 1D features before feeding them into LSTM layers. The primary difference among the models was the number of CNN and LSTM layers. Ultimately, the CNN-LSTM network with two convolutional layers and two LSTM layers was selected for its balance of performance and efficiency. The results showed that the combination of convolution and LSTM methods outperformed standalone LSTM and CNN models.
Furthermore, Marjani et al. (2024b) combined CNN with bidirectional LSTM (BiLSTM) for a study in Australia, using wildfire driving factors from the previous four days and wildfire masks to predict next-day wildfire spread. The method first employed four CNN layers to extract spatial features, which were then flattened and input into a three-layer BiLSTM. The model produced binary outputs through dilated convolution, dense layers, and sigmoid activation functions. Comparative experiments demonstrated that this approach outperformed LSTM and Conv-LSTM methods, and the study also confirmed that using longer time steps improved prediction capabilities. In another study, Li et al. (2021c) not only introduced a CNN before the LSTM but also incorporated an attention mechanism within the LSTM to assign different weights to the hidden states, reducing information loss and improving prediction accuracy. Applied in Portugal’s Montesinho Natural Park, the model used various meteorological factors from the FWI, historical burned areas, and time and location data to predict future wildfire risk. Comparisons with SVM, XGBoost, neural networks, RNN, and basic LSTM showed that deep learning models significantly outperformed machine learning models, with LSTM surpassing RNN and the CNN-attention LSTM performing best overall.
@@ 6.4.2 Coupled Spatiotemporal Modeling Methods
Finally, Eddin et al. (2023) designed two separate convolutional branches to process dynamic variables (those changing over time) and static variables (those remaining constant throughout the study period). To address the causal effects of static variables on dynamic ones, they incorporated a location-aware adaptive normalization layer that adjusts the activation maps in the dynamic branch based on static features. They also employed sinusoidal encoding of Julian days to provide the model with explicit temporal information. Their experiments on the FireCube dataset (Prapas et al., 2022; Kondylatos et al., 2023), which covers parts of the Mediterranean, demonstrated that the proposed model’s predictions of wildfire spread at time $t+1$ outperformed baseline models, including LSTM, ConvLSTM, RF, and XGBoost. In a similar vein, Dong et al. (2024) proposed an architecture that models dynamic and static variables separately during the encoding phase, leading to improved accuracy in predicting fire radiative power.
@@ 6.4.3 Spatially Enhanced Coupled Spatiotemporal Modeling Methods
In addition, there are studies that combine convolutional layers with ConvLSTM to further emphasize spatial feature extraction. Similar to the separated spatial and temporal sequence wildfire risk prediction models, the spatially enhanced coupled spatiotemporal wildfire risk prediction method, as shown in Fig. 6, first extracts spatial features using methods such as convolution operations. Dimensionality reduction is then applied before feeding the data into a temporal model, such as ConvLSTM. The output of the temporal model can then be processed by a convolutional block followed by a sigmoid mapping, or directly mapped using a sigmoid function, depending on the specific design. Compared to the separated spatial and temporal modeling approach, this method tends to preserve a certain level of spatial information throughout the modeling process, which may help retain finer-grained spatial patterns in the final prediction.
Figure 6: A typical spatial feature-enhanced coupled spatiotemporal wildfire risk prediction model diagram.
@@
Similarly, He et al. (2024) developed a next-day wildfire risk prediction model for eastern China by combining CNN, ConvLSTM, and Vision Transformer models, also employing a branch structure. In this model, eight CNN modules were used to model eight meteorological variables, while another CNN module modeled terrain, vegetation, and anthropogenic factors from the last day of each time step. The sequence feature maps generated by the temporal meteorological variables were input into LSTM, and its output was merged with the other variables’ CNN outputs. Channel attention, spatial attention, and CNN were used to extract higher-order features. To emphasize the high-resolution structural features of vegetation and scattered small settlements, Landsat data were used to generate 30-meter monthly high-resolution NDVI products, and a Vision Transformer was applied to extract these structural features. The flattened outputs from the Vision Transformer and CNN were concatenated and classified using an ANN binary classifier. Ablation studies demonstrated the effectiveness of incorporating attention mechanisms, high-resolution spatial structure features, and Vision Transformer. Sensitivity analysis also confirmed that meteorological factors were the most important among wildfire drivers.
@@
Finally, Chen et al. (2024a) designed an end-to-end encoder-decoder network structure similar to that of Bhowmik et al. (2023) and Yoon and Voulgaris (2022), combining GCN and ConvLSTM. The GCN component was introduced to model the long-distance spatial dependencies required for global wildfire risk prediction, while mitigating the impact of missing data from ocean regions. Explicitly modeling information flow through the graph edges improved the interpretability of the prediction results. Specifically, the Earth’s land surface pixels, as simulated by the JULES-INFERNO model, were treated as graph nodes, and the correlation coefficients between nodes were used to establish the graph’s edges. Spatial dependencies were modeled using a spatial graph convolution model, followed by LSTM for long-term temporal dependencies. The model was trained and validated using 30 years of simulated data from 1961 to 1990 in the JULES-INFERNO dataset, and its performance was compared against LSTM, ConvLSTM, and convolutional autoencoder with LSTM. The proposed model demonstrated superior performance, with ConvLSTM being the best among the comparative models. Additionally, the study used the Louvain method to identify potential wildfire communities, and integrated gradient analysis revealed that short-term (one-month) wildfire risk is primarily driven by lightning, while long-term (11-month) risk is more influenced by temperature, followed by lightning. Node importance analysis also showed that even geographically distant locations could have strong interactions regarding wildfire risk. Similarly, Zhao et al. (2024) aimed to improve model interpretability by combining causal GNN with LSTM, comparing its performance with LSTM and GRU in predicting future wildfire risk.
This section reviews recent advancements in deep learning methods applied to wildfire risk prediction, emphasizing their evolution from temporal and spatial modeling to spatiotemporal sequence modeling and the integration of emerging technologies. Due to their hierarchical structures and flexible architectures (LeCun et al., 2015), deep learning models excel at capturing nonlinear relationships, enabling them to model the complex interactions and latent patterns within high-dimensional spatiotemporal data sourced from meteorological variables, vegetation indices, topographical features, and socioeconomic factors. Unlike traditional statistical models that rely on explicit feature engineering and linear assumptions, deep learning models uncover intricate dependencies that are difficult to detect using linear or shallow frameworks. This capability is particularly crucial in wildfire risk prediction, where the interdependencies among driving factors often exhibit nonlinear behaviors that vary dynamically across both temporal and spatial scales (Jain et al., 2020).
The evolution of deep learning models in this domain demonstrates significant strides in both predictive accuracy and computational efficiency (Cheng et al., 2024). Early applications relied on RNNs and LSTM networks to process wildfire time series data, while CNNs and GCNs were used to extract spatial features. Recent innovations have merged these approaches into hybrid spatiotemporal models, enabling simultaneous analysis of spatial and temporal dependencies. These advancements have been accompanied by the introduction of unique design elements tailored to address specific challenges in wildfire risk prediction. For example, dynamic-static feature separation enables models to distinguish between temporally varying inputs, such as meteorological variables, and static attributes, such as topography and land cover, improving prediction efficiency (Eddin et al., 2023; Shadrin et al., 2024; Jiang et al., 2023; Dong et al., 2024). Explicit modeling of temporal and spatial dependencies, often through attention mechanisms and spatiotemporal encoding, allows these models to capture complex interactions across time steps or spatially distant regions, thereby enhancing their ability to predict wildfire propagation over large areas (Ji et al., 2024b). Additionally, graph-based representations of geographical regions, where nodes correspond to areas and edges represent interactions such as wind patterns or vegetation continuity, provide a new framework for modeling long-range dependencies and spatial interactions inherent in wildfire dynamics (Jiang et al., 2022; Chen et al., 2022, 2024a).
7 Promising Deep Learning Methods and Data
In recent years, the model architectures and the types and quantity of driving features employed in wildfire risk prediction have largely stabilized. From a modeling perspective, given the complex nonlinear and spatiotemporal relationships between various driving factors and wildfire occurrence and spread, a key trend in deep learning-based wildfire prediction is enhancing the capability to model long-range spatiotemporal dependencies. This aims to capture subtle and weak correlations effectively. More specifically, wildfire risk prediction has evolved from purely temporal sequence forecasting to spatiotemporal sequence prediction and spatially enhanced spatiotemporal modeling. These advancements focus on increasing both the temporal and spatial scales of the model, enabling the incorporation of higher-resolution data over longer time sequences while expanding the dataset size. Current research has explored wildfire risk prediction models based on RNNs, CNNs, transformers, and GNNs.
However, compared to more recent models such as Mamba (Gu and Dao, 2023), these approaches still suffer from inefficiencies or limited capability in modeling large-scale spatiotemporal relationships. Additionally, many of these models rely on multi-branch networks that primarily use image-based or node-based representations to capture long-range semantic dependencies, without exploring the potential efficiency gains offered by multimodal approaches, such as incorporating textual descriptions. Furthermore, the applicability of cross-disciplinary solutions in addressing the inherent challenges of existing wildfire risk prediction models—such as data dependency, poor interpretability, difficulty in uncertainty quantification, and challenges in handling stochasticity, multi-source heterogeneous data, and complex factor interactions—remains an open question.
From a data perspective, large-scale wildfire risk prediction studies primarily prioritize data availability, commonality, and ease of processing. This has resulted in an over-reliance on meteorological data from various sources, particularly ground station measurements, ERA-5 reanalysis data, and multiple MODIS-derived products. While MODIS provides high temporal resolution and long-term large-scale observations, its passive optical remote sensing capabilities are inherently limited in penetrating vegetation canopies. Consequently, MODIS-based assessments of fuel conditions are often imprecise, and the sensor’s inability to capture data under all-weather, all-day conditions significantly constrains the applicability of wildfire risk prediction models trained on such data. A key question that remains is whether alternative data sources can improve the accuracy of wildfire risk prediction, particularly given the fundamental reality that data acquisition capabilities in Earth sciences far exceed current data processing capacities (Karniadakis et al., 2021).
In response to these challenges, this section systematically reviews recent advancements in deep learning methods—particularly those employing multimodal approaches—for segmentation, classification, and prediction tasks. Additionally, it provides a brief overview of the Mamba model, a parameter-efficient deep learning framework, to facilitate the development of computationally efficient wildfire risk prediction models in future research. Given the similarities between meteorological forecasting and wildfire risk prediction—both of which involve multiple driving factors, nonlinear coupling, and dynamic spatiotemporal modeling—this section also summarizes recent developments in meteorological forecasting models. The goal is to explore whether cross-disciplinary methodologies can enhance deep learning-based wildfire risk prediction by introducing physical constraints, improving uncertainty quantification, and incorporating lightweight model designs to enhance reliability and accuracy. Furthermore, this section discusses underutilized remote sensing data types in wildfire risk prediction, particularly microwave and LiDAR data, which offer penetrative capabilities, as well as high-spatial-resolution nighttime light data, which can enable dynamic nighttime wildfire monitoring. These discussions aim to provide a broader data reference for future wildfire risk research.
7.1 Promising Deep Learning Models
@@ 7.1.1 Multimodal Models
Decoder fusion extracts visual and language features separately using different encoders, such as CNNs for visual data and LSTMs for language data, and designs specific fusion modules to combine these features at the decoder stage. For example, Jain and Gandhi (2022) utilized CNNs to extract multi-stage image features and LSTMs for text feature extraction, integrating these features through a hierarchical cross-modal aggregation module. In another approach, Wang et al. (2022b) designed a knowledge transfer method to adapt image-level knowledge from Contrastive Language-Image Pretraining (CLIP) into RIS for pixel-level segmentation tasks, implementing a Neck module to fuse outputs from both the text and image encoders. Similarly, Yan et al. (2023) introduced a Fusion Neck module to combine image and text encoding information, alongside a multi-query generator module to aid segmentation.
@@
In summary, multimodal tasks offer greater flexibility and intuitive interaction by incorporating natural language to guide the model, compared to unimodal tasks. Multimodal information provides additional contextual data that unimodal models often lack, allowing the language encoder to construct a relation graph for numerical geospatial relations (Yuan et al., 2024). Therefore, multimodal models, especially RIS, hold significant potential in wildfire risk prediction by capturing subtle spatiotemporal semantic dependencies that are difficult to model using existing methods.
7.1.2 Efficient Deep Learning Models